# A theoretical framework for self-supervised contrastive learning for continuous dependent data

### v1

@@ Abstract
Self-supervised learning (SSL) has emerged as a powerful approach to learning representations, particularly in the field of computer vision. However, its application to dependent data, such as temporal and spatio-temporal domains, remains underexplored. Besides, traditional contrastive SSL methods often assume semantic independence between samples, which does not hold for dependent data exhibiting complex correlations. We propose a novel theoretical framework for contrastive SSL tailored to continuous dependent data, which allows the nearest samples to be semantically close to each other. In particular, we propose two possible ground truth similarity measures between objects — hard and soft closeness. Under it, we derive an analytical form for the estimated similarity matrix that accommodates both types of closeness between samples, thereby introducing dependency-aware loss functions. We validate our approach, Dependent TS2Vec, on temporal and spatio-temporal downstream problems. Given the dependency patterns presented in the data, our approach surpasses modern ones for dependent data, highlighting the effectiveness of our theoretically grounded loss functions for SSL in capturing spatio-temporal dependencies. Specifically, we outperform TS2Vec on the standard UEA and UCR benchmarks, with accuracy improvements of $4.17$% and $2.08$%, respectively. Furthermore, on the drought classification task, which involves complex spatio-temporal patterns, our method achieves a $7$% higher ROC-AUC score.
@@ III-C1 Ground truth similarity matrices
Hard dependency (moving average). Here, neighbor observations $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ with close $i$ and $j$ such that $|i-j|=1$ are dependent. Thus, closely lying samples can be treated as positive ones. Then, we can introduce the following model for the matrix $\mathbf{G}$:
@@ III-C2 Estimated similarity matrix
$\widehat{\mathbf{G}}_{d,\mathcal{R}}=\operatorname*{arg\,min}_{\mathbf{G}\in\mathcal{G}}\mathrm{Tr}(\mathbf{D}\mathbf{G})+\mathcal{R}(\mathbf{G}),$ (6)
[…]
@@
$\frac{\partial{\mathcal{L}}}{\partial{\mathbf{G}_{ij}}}=\begin{cases}d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))+\tau\ln(g_{ij})+\alpha_{i},&j>i,\\ d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))+\tau\ln(g_{ij})+\beta_{j},&j<i,\\ \end{cases}$
@@ IV-B4 SSL architecture
We adopt the TS2Vec encoder [13], a multi-layer temporal convolutional network that generates hierarchical representations. Our temporal loss is applied to the resulting embeddings during self-supervised training, see details in Appendix D-A.
@@ IV-B5 Results
In Table I we present the classification results on UCR and UEA archives using average accuracy and average rank metrics. Our proposed method surpasses previous approaches on both UCR and UEA archives. In particular we achieve the highest mean accuracy on both benchmarks – $84.64$% on UCR and $72.89$% on UEA – outperforming TS2Vec by $2.08$% and $4.17$%, and SoftCL by $1.06$% and $1.20$%, respectively.
TABLE I: Comparison of average classification accuracy ()($\uparrow$)( ↑ ) and average rank ()($\downarrow$)( ↓ ) on UCR and UEA archives. Best results are marked in bold, second one are underlined.
@@
TABLE II: Mean rank ()($\downarrow$)( ↓ ) by dataset type across methods on UCR and UEA benchmarks.
@@ IV-C1 Downstream problems description
Instead of solving the regression problem, i.e., predicting PDSI directly, we consider a binary classification task: whether it is a drought or not. To convert the regression task to a classification problem, we considered a case as a drought if the PDSI is less than $-2$ following [38].
@@ IV-C3 Data
  • PDSI index collected for four different regions — Madhya Pradesh (India), Missouri (USA), Northern Kazakhstan, and Eastern Europe — to estimate the drought prediction quality. The data covers a time from $1958$ to $2022$ with a monthly resolution. We utilized the TerraClimate Monthly dataset  [^1] https://developers.google.com/earth-engine/datasets/catalog/IDAHO_EPSCOR_TERRACLIMATE to obtain the PDSI data for chosen regions. Their brief characteristics are summarized in Appendix C.
@@ IV-C4 SSL model architecture
In Appendix D-B we provide the implementation details for the model.
@@ IV-C5 Results
TABLE III: ROC-AUC ()($\uparrow$)( ↑ ) values for drought prediction for different forecasting horizons and climatic regions
@@ V Conclusions
To evaluate our theoretical framework in practice, we made an extensive experimental evaluation. Our method consistently outperformed state-of-the-art baselines such as TS2Vec and SoftCL. In time series classification tasks on the UCR and UEA benchmarks, DepTS2Vec achieved accuracy gains of $2.08$% and $4.17$%, respectively, over TS2Vec. For spatio-temporal downstream problems such as drought classification, our approach delivered a $7$% improvement in ROC-AUC, and it also achieved the lowest RMSE in temperature forecasting, outperforming both TS2Vec and SoftCL.
@@ Appendix A Proof of Theorem 1 from Section III
$\frac{\partial{\mathcal{L}}}{\partial{g_{ij}}}=\begin{cases}d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))+\tau\ln(g_{ij})+\alpha_{i},&j>i,\\ d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))+\tau\ln(g_{ij})+\beta_{j},&j<i,\\ \end{cases}$
@@
$j>i:\begin{cases}d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))+\tau\ln(g_{ij})+\alpha_{i}=0\\ \sum_{j=i+1}^{N}g_{ij}-1=0,\\ \end{cases}$
$j<i:\begin{cases}d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))+\tau\ln(g_{ij})+\beta_{j}=0\\ \sum_{i=j+1}^{N}g_{ij}-1=0,\\ \end{cases}$
@@
$g_{ij}=\begin{cases}\frac{e^{\frac{-1}{\tau}d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))}}{\sum_{j=i+1}^{N}e^{\frac{-1}{\tau}d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))}},&j>i\\ \frac{e^{\frac{-1}{\tau}d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))}}{\sum_{i=j+1}^{N}e^{\frac{-1}{\tau}d(f_{\theta}(\mathbf{x}_{j}),f_{\theta}(\mathbf{x}_{i}))}},&j<i\\ \end{cases}$
@@
Appendix B Exponential semantic relations between samples
Below we show another way how to model correlations between samples. In particular we can consider the exponential closeness between elements which can be formalized in the following way:
$g_{ij}=\begin{cases}\exp\left(-\frac{(i-j)^{2}}{k}\right),&\text{if}\ i\neq j,\\ 0,&\text{otherwise}.\end{cases}$ (7)
with $k>0$. Before we can formulate the corresponding optimization problem, we need to find the constraint for this variant of a matrix.
Lemma B.1.
Let $\mathbf{G}$ be a similarity matrix defined by  (7). Then
$\sum_{i\neq j}\ln(g_{ij})=\frac{N^{2}(1-N^{2})}{6k}.$ (8)
where $k>0$.
Proof.
First, let us note that $\sum_{i\neq j}\ln(g_{ij})=\sum_{i,j}\ln\left(\exp\left(-\frac{(i-j)^{2}}{k}\right)\right).$
Now, we should calculate the right side of the expression.
Consider one element under the logarithm
$a_{ij}=\exp\left(-\frac{(i-j)^{2}}{k}\right)\Longrightarrow$
$\ln(a_{ij})$ $=-\frac{(i-j)^{2}}{k}=-\frac{i^{2}-2ij+j^{2}}{k}$ (9)
$=-\frac{i^{2}}{k}+\frac{2ij}{k}-\frac{j^{2}}{k}.$
Summing all elements across the columns, we obtain:
$\sum_{j=1}^{N}\ln(a_{ij})=-\frac{1}{k}\sum_{j=1}^{N}j^{2}+\frac{2i}{k}\sum_{j=1}^{N}j-\frac{Ni^{2}}{k}\\ =-\frac{N(N+1)(2N+1)}{6k}+i\frac{N(N+1)}{k}-\frac{Ni^{2}}{k},\$
Similarly, the sum of the last two terms, which are depending on the row indexes, are the following:
$\sum_{i=1}^{N}i\frac{N(N+1)}{k}$ $=\frac{N^{2}(N+1)^{2}}{2k},$
$\sum_{i=1}^{N}\frac{Ni^{2}}{k}$ $=\frac{N^{2}(N+1)(2N+1)}{6k},$
Thus, the final expression for the constraint is
$\sum_{i,j}\ln(a_{ij})=\frac{N^{2}(1-N^{2})}{6k}.$
Similarly to the nearby relation we also proposed the constraint from Lemma B.1 to account for specific properties of the $\mathbf{G}$. Hence, the optimization problem for the exponential relation case is the following:
$\widehat{\mathbf{G}}_{d,\mathcal{R}}$ $=\operatorname*{arg\,min}_{\mathbf{G}\in\mathcal{G}}\mathrm{Tr}(\mathbf{D}\mathbf{G})+\mathcal{R}(\mathbf{G}),$ (10)
[…]
Solving optimization problem 10 leads to the following solution:
Theorem 3.
Solving the optimization problem (10) using contsraint from (B.1) without regularizer gives the following estimation of the similarity matrix:
$(\widehat{g}_{d})_{ij}=\begin{cases}\frac{\exp\left(\frac{1}{N^{2}-N}(C+\sum_{i\neq j}\ln\left(d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))\right)\right)}{d(f_{\theta}(\mathbf{x}_{i}),f_{\theta}(\mathbf{x}_{j}))},&i\neq j,\\ 0,&i=j,\end{cases}$
where
$C=\frac{N^{2}(1-N^{2})}{6k},k>0.$
Note: constant $C$ was obtained in Lemma B.1.
Proof.
The proof is similar to Theorem 1. ∎
For the exponential relation we don’t have any regularizer. However non-linear contstraint delivers the minimum.
Theorem 4.
The estimated similarity matrix from Theorem 3 is a unique minimum of the optimization problem 10.
Proof.
The proof is similar to Theorem 2. ∎
Due to the numerical instability of the estimated similarity matrix from the Theorem 3, we don’t include it in our experimental results.
Appendix C Drought data statistics
Table IV provides accumulated statistics on time coverage and percentage of the droughts for four considered regions. We see that the range of data is similar over all regions, while the share of droughts varies from $25\%$ to $31\%$, demonstrating diverse conditions.
TABLE IV: Region’s statistics
[…]
months PDSI $\geq-2$ PDSI $\leq-2$
Missouri, USA 754 74.91 25.09
Madhya Pradesh, India 754 70.66 29.34
Europe 742 66.28 33.72
Northern Kazakhstan 742 68.70 31.30
Appendix D Technical details for DepTS2Vec
In this section, we describe technical implementation details for our DepTS2Vec approach. Since temporal and spatio-temporal data require different settings, we divide our section into two parts, devoted to temporal and spatio-temporal data, correspondingly.
D-A DepTS2Vec for temporal data
@@
D-B DepTS2Vec for spatio-temporal data

### v2

@@ Abstract
Self-supervised learning (SSL) has emerged as a powerful approach to learning representations, particularly in the field of computer vision. However, its application to dependent data, such as temporal and spatio-temporal domains, remains underexplored. Besides, traditional contrastive SSL methods often assume semantic independence between samples, which does not hold for dependent data exhibiting complex correlations. We propose a novel theoretical framework for contrastive SSL tailored to continuous dependent data, which allows the nearest samples to be semantically close to each other. In particular, we propose two possible ground truth similarity measures between objects — hard and soft closeness. Under it, we derive an analytical form for the estimated similarity matrix that accommodates both types of closeness between samples, thereby introducing dependency-aware loss functions. We validate our approach, Dependent TS2Vec, on temporal and spatio-temporal downstream problems. Given the dependency patterns presented in the data, our approach surpasses modern ones for dependent data, highlighting the effectiveness of our theoretically grounded loss functions for SSL in capturing spatio-temporal dependencies. Specifically, we outperform TS2Vec on the standard UEA and UCR benchmarks, with accuracy improvements of 4.174.17% and 2.082.08%, respectively. Furthermore, on the drought classification task, which involves complex spatio-temporal patterns, our method achieves a 77% higher ROC-AUC score.
@@ III-C1 Ground truth similarity matrices
Hard dependency (moving average). Here, neighbor observations $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ with close $i$ and $j$ such that |ij|=1|i-j|=1 are dependent. Thus, closely lying samples can be treated as positive ones. Then, we can introduce the following model for the matrix $\mathbf{G}$:
@@ III-C2 Estimated similarity matrix
$\textrm{s.t.}\sum_{j=i+1}^{N}g_{ij}$ $=1\qquad\text{for any\,\,}i\in\overline{\{1,N\}},$
@@
[…]
@@ IV-B4 SSL architecture
We adopt the TS2Vec encoder [13], a multi-layer temporal convolutional network that generates hierarchical representations. Our temporal loss is applied to the resulting embeddings during self-supervised training, see details in Appendix A-A.
@@ IV-B5 Results
In Table I we present the classification results on UCR and UEA archives using average accuracy and average rank metrics. Our proposed method surpasses previous approaches on both UCR and UEA archives. In particular we achieve the highest mean accuracy on both benchmarks – 84.6484.64% on UCR and 72.8972.89% on UEA – outperforming TS2Vec by 2.082.08% and 4.174.17%, and SoftCL by 1.061.06% and 1.201.20%, respectively.
TABLE I: Comparison of average classification accuracy $(\uparrow)$ and average rank $(\downarrow)$ on UCR and UEA archives. Best results are marked in bold, second one are underlined.
@@
TABLE II: Mean rank $(\downarrow)$ by dataset type across methods on UCR and UEA benchmarks.
@@ IV-C1 Downstream problems description
Instead of solving the regression problem, i.e., predicting PDSI directly, we consider a binary classification task: whether it is a drought or not. To convert the regression task to a classification problem, we considered a case as a drought if the PDSI is less than 2-2 following [38].
@@ IV-C3 Data
  • PDSI index collected for four different regions — Madhya Pradesh (India), Missouri (USA), Northern Kazakhstan, and Eastern Europe — to estimate the drought prediction quality. The data covers a time from $1958$ to $2022$ with a monthly resolution. We utilized the TerraClimate Monthly dataset  [^1] https://developers.google.com/earth-engine/datasets/catalog/IDAHO_EPSCOR_TERRACLIMATE to obtain the PDSI data for chosen regions.
@@ IV-C4 SSL model architecture
In Appendix A-B we provide the implementation details for the model.
@@ IV-C5 Results
TABLE III: ROC-AUC $(\uparrow)$ values for drought prediction for different forecasting horizons and climatic regions
@@ V Conclusions
To evaluate our theoretical framework in practice, we made an extensive experimental evaluation. Our method consistently outperformed state-of-the-art baselines such as TS2Vec and SoftCL. In time series classification tasks on the UCR and UEA benchmarks, DepTS2Vec achieved accuracy gains of 2.082.08% and 4.174.17%, respectively, over TS2Vec. For spatio-temporal downstream problems such as drought classification, our approach delivered a 77% improvement in ROC-AUC, and it also achieved the lowest RMSE in temperature forecasting, outperforming both TS2Vec and SoftCL.
@@ Appendix A Proof of Theorem 1 from Section III
[…]
@@
[…]
@@
[…]
@@
A-A DepTS2Vec for temporal data
@@
A-B DepTS2Vec for spatio-temporal data